OpenStudy (anonymous):

Suppose $500 is deposited into an account paying interest at a rate of 7%, continuously compounded. Find a formula for the value of the account at time t. What is the value of the account after 3 years?

OpenStudy (anonymous):

compounded pay twice a years A= 500(1+0.07/2)^(2*3)

OpenStudy (anonymous):

$614.28 If I deserve medal then give me one.

OpenStudy (dumbcow):

It is being compounded continuously B = 500*e^.07*3 B = 500*e^.21 B = 616.84